A snow plough with a mass of m=8 metric tonnes (1 metric tonne=1,000kg) has an initial speed v1 =14 m/s is clearing snow. The retarding force the snow produces on the snow plough satisfies the force-displacement relationship, Fsnow = 900+6000x2 where Fsnow is in Newtons and x is in metres. Each wheel have a radius of 0.5m and a mass moment of inertia of 3.75 kgm2. Assuming the wheels do not slip and friction can be ignored what is the speed of the snow plough after it travels a distance s=3m?
Answer #1

Given that

a mass of m=8 metric tonnes(1 metric tonne=1,000 kg)

v1=14 m/s

F snow=900+6000 x^2

a distance s=3 m

radius of 0.5 m

Answer force on the snow Plough, f= 900+ 6000 ..ale = fim = 18000 [ 900+booth vor s as de- acteleration here? now, as-vd y